OpenStudy (anonymous):

how? :( an unknown gas contains 85.6% c and 14.4% h. at 0.455 atm and 425 k, the gas has a density of 0.549 g/l. what is the molecular formula for the gas?

OpenStudy (anonymous):

@thomaster :(

OpenStudy (anonymous):

\[PV=NRT\] N=mass/molar mass=M/Mm \[PV=M/Mm*R*T\] Density=mass/volume=M/V \[P=D*R*T/Mm\] Solve for Mm \[Mm=DRT/P\]=42.08g grams of C=Grams total*percent carbon \[42.8*0.856=36.02\] moles carbon=grams carbon/molar mass 36.02/12=3 Do the same thing for H \[42.8*0.144=6.06\] 6.06/1.01=6 The formula is C3H6

OpenStudy (anonymous):

n / V = P / RT = (0.455 atm) / ((0.08205746 L atm/K mol) x (425 K)) = 0.01305 mol/L (0.549 g/L) / (0.01305 mol/L) = 42.069 g/mol Supposing the given percentages are percents by mass: (42.069 g/mol) x (0.856) / (12.01078 g C/mol) = 2.998 (42.069 g/mol) x (0.144) / (1.007947 g H/mol) = 6.010 Round to the nearest while numbers to find the molecular formula: C3H6